OpenStudy (anonymous):

@jim_thompson5910 need help before 11:59 tonight please help 3. The length of a rectangle is 7 mm longer than its width. Its perimeter is more than 62 mm. Let w equal the width of the rectangle. a. Write an expression for the length in terms of the width. b. Use these expressions to write an inequality based on the given information. c. Solve the inequality, clearly indicating the width of the rectangle.

OpenStudy (academicgurusinc):

L = length W= width P= perimeter L= w + 7 since P=2L + 2W = 2(L+W) and P>62 2(W + 7) + 2w > 62 2W + 14 + 2w > 62 4W + 14 > 62 4W > 62 - 14 4W > 48 W > 48/4 W > 12 Therefore, the width is greater than than 12 mm

OpenStudy (anonymous):

Length = L , Width = W , (L=W+7) since the length is longer than the width by 7 mm Perimeter = 2L+2W , substituting from previous equation , Perimeter =2W +14+2W=4W+14. perimeter>62 , Then 4W+14>62 , 4W>48 ,W>12 , Same With Length (W=L-7) You Will Get 4L-14>62 L>19

OpenStudy (academicgurusinc):

We hope this helps. Regards, AcademicGurusInc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc)